# Generating Symbolic Music from Natural Language Prompts using an LLM-Enhanced Dataset

### v2

@@ IV Methods
V Experiments and Results
V-A Baseline Models
We compare our method to an existing work for symbolic music with ABC notations with the same prompts that we randomly pick from our test set. BART-base [10] investigate the efficacy of pre-trained checkpoints in text-to-music generation. We generate music with BART-base model in ABC notation using its Hugging Face API [32]. For a fair comparison, we convert the generated ABC notation to MIDI using Melobytes, an online software which supports multitrack conversion [33]. We intended to compare our proposed model to MuseCoco[3]. However, not all free-form natural language captions in our test set can conform to their template.
V-B Objective Evaluation
Following [1, 27, 26], we use pitch class entropy, scale consistency and groove consistency to evaluate generated music. We randomly pick 100 samples from the test set to compute these metrics, and we consider a closer value to that of the ground truth better. We report in Table IV the pitch class entropy, scale consistency and groove consistency of MST-Tags-Small, MST-Tags, MST-Text and the ground truth. We consider the generated music better if the corresponding value is closer to that of the ground truth. We observe that MST-Tags-Small and MST-Tags lead to values significantly different from those of the ground truth in terms of pitch class entropy and scale consistency.
Further, MST-Text is able to generate music that is statistically close to the ground truth in terms of all three metrics. However, we note that none of these objective evaluation metrics measures the tags/text-to-music adherence, and we will examine this in the subjective listening test.
V-C Subjective Evaluation
@@
We report the subjective evaluation results in Table III. When contrasting MST-Tags-Small with MST-Tags, we observe that MST-Tags achieves better performance in coherence and arrangement, but we see a decrease in adherence, possibly due to the incorporation of some auto-generated tags. This comparison illustrates the trade-off between employing a smaller, high-quality dataset (MetaScore-Genre) versus a larger yet noisy dataset (MetaScore-Plus). However, comparing the overall quality score of MST-Tags-Small and MST-Tags, we see that training with a larger dataset leads to an increase in the overall quality of music generation.
@@
For text-conditioned music generation, MST-Text outperforms the BART-based[10] approach, in terms of coherence, arrangement, adherence, and overall quality. We observe that text-conditioned system MST-Text has a lower adherence against MST-Tags-Small and MST-Tags. However, Table II shows our generated captions keep all the information with in-context learning. This implies that text-to-music generation is a more challenging task than tag-to-music generation because additional effort is required to interpret free-form text for music generation.
@@
Overall, the tag-conditioned music generation systems, MST-Tags and MST-Tags-Small, perform very well, suggesting the high quality of our constructed dataset. In addition, MST-Text achieves the highest overall quality, which shows that our text-conditioned system performs similarly well with our tag-conditioned system. This showcases the effectiveness of our approach that leverages a large language model to generate natural language captions for end-to-end training of text-to-music generation models.
VI Conclusion
@@
In this work, we have presented a new music generation model that learns to generate symbolic music from free-form text, allowing controls over instruments, genre, composer, complexity and other free-form music descriptors. In addition, the LLM-generated pseudo captions contain information provided in free-form user-annotated tags, which can poses a challenge to systems that adopt a predefined set of tags, including our tag-to-music model and the two-stage approach proposed in [3]. Our objective and subjective evaluation results show the effectiveness of the proposed tags-to-music and text-to-music models, both outperforming a baseline text-to-music model [10]. In addition, the proposed text-to-symbolic music generation model trained with LLM-generated pseudo captions achieves competitive performance against the proposed tags-to-music model trained using only the ground truth tags.
@@
Finally, we would like to point out that, as a generative model trained on copyrighted material, our proposed model has the potential to generate samples that could lead to copyright infringement. However, when used properly and with cautions, a text-to-music generation system can also make a positive impact to the society by enabling new opportunities and interfaces for music creation, as demonstrated in [28]. Given the editable nature of symbolic music, we hope our proposed text-to-symbolic music models to open up new pathways towards human-AI music co-creation.

### v3

@@ 6.1 Baselines
6.2 Objective Evaluations
Following [15, 24, 25], we assess the quality of generated music using pitch class entropy, scale consistency, and groove consistency, where values closer to the ground truth indicate better performance. To ensure a fair comparison, we randomly sampled 100 conditions from the MST test set and generated corresponding music for evaluation. As reported in Table 4, we find that MST-Text most closely matches the ground truth in terms of pitch class entropy and scale consistency, while MST-Tags-Small and MST-Text perform similarly on groove consistency. Additionally, our proposed MST-Text outperforms the BART-based[13] model across all three metrics.
Model size | Training samples | Coherence | Arrangement | Adherence | Overall quality | |
MST-Tags-Small | 87.36M | 150K | 3.87 0.36 | 3.98 0.38 | 3.86 0.38 | 3.57 0.37 |
MST-Tags | 87.36M | 901K | 4.01 0.37 | 4.06 0.39 | 3.60 0.49 | 3.66 0.45 |
BART-based[13] | 139M | 283K | 3.86 0.30 | 3.63 0.39 | 2.81 0.50 | 3.29 0.42 |
MST-Text | 87.44M | 560K | 3.93 0.28 | 3.88 0.33 | 3.35 0.44 | 3.69 0.33 |
Model size | CLAP Score | Coherence(%) | Arrangement(%) | Adherence(%) | Overall quality(%) | |||||||||||
M | T | M+T | M | T | M+T | M | T | M+T | M | T | M+T | M | T | M+T | ||
Text2MIDI [1] | 159M | 0.23 | 0.20 | 0.22 | 0 | 40 | 20 | 40 | 50 | 45 | 40 | 80 | 60 | 40 | 60 | 50 |
MST-Text | 87.44M | 0.36 | 0.13 | 0.24 | 100 | 60 | 80 | 60 | 50 | 55 | 60 | 20 | 40 | 60 | 40 | 50 |
6.3 Subjective Evaluation
@@
We report the subjective evaluation results in Table 5. When contrasting MST-Tags-Small with MST-Tags, we observe that MST-Tags achieves better performance in coherence and arrangement, but we see a decrease in adherence, possibly due to the incorporation of some auto-generated tags. This comparison illustrates the trade-off between employing a smaller, high-quality dataset (MetaScore-Genre) versus a larger yet noisy dataset (MetaScore-Plus). However, comparing the overall quality score of MST-Tags-Small and MST-Tags, we see that training with a larger dataset leads to an increase in the overall quality of music generation.
@@
For text-conditioned music generation, MST-Text outperforms the BART-based[13] approach, in terms of coherence, arrangement, adherence, and overall quality. We observe that the text-conditioned system MST-Text has a lower adherence against MST-Tags-Small and MST-Tags. This implies that text-to-music generation is a more challenging task than tag-to-music generation as a text-to-music generation system needs to learn to interpret the free-form text inputs.
@@
Overall, the tag-conditioned systems, MST-Tags and MST-Tags-Small, demonstrate strong performance across multiple dimensions, including coherence, arrangement, adherence to prompts, and overall music quality. These results highlight the high quality of our constructed dataset. Notably, MST-Text achieves the highest score in overall quality, indicating that our text-conditioned model performs on par with the tag-conditioned variants. This underscores the effectiveness of our approach, which leverages a large language model to generate natural language captions, enabling end-to-end training for high-quality text-to-music generation.
6.4 Comparison to Text2MIDI
In this section, we compare our proposed MST-Text with a concurrent work Text2MIDI[1] that also supports text-to-symbolic music generation. For a fair comparison, we create three test sets: 1) five text prompts randomly selected from our test set, 2) five text prompts randomly selected from the Text2MIDI test set, and 3) the union of the previous two test sets (i.e., five prompts from each test set).
Objective Evaluation. We compare music quality of our MST-Text and Text2MIDI [1] using pitch class entropy, scale consistency, and groove consistency, and report the results in Table 7. In addition, to assess the alignment between text and symbolic music, we report the average semantic similarity computed with CLAP [17]. We find that MST-Text better matches the ground truth in terms of pitch class entropy and scale consistency, while Text2MIDI better matches the ground truth in terms of groove consistency. Additionally, MST-Text shows better alignment performance when prompts are taken from the MST-Text test set and the joint test set.
Pitch class entropy | Scale consistency | Groove consistency | |
Text2MIDI[1] | 2.44 0.19 | 0.89 0.03 | 0.94 0.01 |
MST-Text | 2.65 0.08 | 0.96 0.02 | 0.92 0.01 |
Ground truth | 2.71 0.07 | 0.96 0.02 | 0.94 0.01 |
Subjective Evaluation. We compare our model MST-Text with Text2MIDI [1] via an A/B test on coherence, arrangement, adherence, and overall quality. Eleven participants (9 with musical experience, including one professional) evaluated 5 prompts from our test set and 5 from the Text2MIDI test set, with results summarized in Table 6.
Our experiments show that MST-Text produces more coherent results and achieves equal or superior arrangement performance compared to Text2MIDI[1]. However, when it comes to adherence, Text2MIDI outperforms MST-Text on the Text2MIDI test set and the joint test sets, while MST-Text performs better on MST-Text test set. Overall, both models deliver comparable quality.
7 Conclusion
@@
In this paper, we have introduced MetaScore, a new publicly available dataset containing rich metadata and LLM-generated captions. We also present a new music generation model that can generate symbolic music from free-form text, allowing controls over instruments, genre, composer, complexity, among other music descriptors. In addition, the LLM-generated pseudo captions contain information provided in free-form user-annotated tags, which can pose a challenge to systems that adopt a predefined set of tags [12]. Our objective and subjective evaluation results show the effectiveness of the proposed tags-to-music and text-to-music models. Our proposed text-to-music model outperforms a baseline text-to-music model [13] and achieves comparable performance with a concurrent work[1]. In addition, the proposed text-to-symbolic music generation model trained with LLM-generated pseudo captions achieves competitive performance against the proposed tags-to-music model trained using only the ground truth tags.
8 Ethics Statement
@@
We note that MetaScore contains many copyrighted contents. This raises concerns of potential misuse of this dataset that can lead to severe copyright infringements. To minimize such risks, we will release those in public domain only and those not in public domain will be shared upon request and only use for research pruporse. Further, music generation systems built upon the MetaScore dataset may infringe the copyright held by the content creators, and thus we must be careful about adopting these systems in commercial applications. However, we would like to point out that when used properly and with caution, a text-to-music generation system can also make a positive impact to society by enabling new opportunities and interfaces for music creation, as demonstrated in [26]. Given the editable nature of symbolic music, we hope our proposed text-to-symbolic music models will open up new pathways towards human-AI music co-creation.